September 1, 2021 Calorimeter

From GlueXWiki
Jump to: navigation, search

'Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. CCAL HDCCAL log book
  4. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
  5. Photon Reconstruction Task Force

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term CALWG Items
  2. Long term items
  3. Calorimeter work packages

Tentative Agenda

  1. Announcements
  2. Run Preparations
    1. FCAL
    2. BCAL
    3. CCAL
  3. Calibration
  4. Monitoring
    1. BCAL LED
    2. BCAL (offline)
    3. FCAL (offline)
  5. Reconstruction
  6. Simulations
  7. Action Items
  8. Any other business


Attendees: Sean, Susan, Tolga, Rebecca, Varun, Malte, Matt, Jon, Karthik, Beni, Mark, Elton, Zisis

  1. Announcements: none.
  2. Run Preparations
    1. FCAL: a massive amount of integrated work has gone into the FCAL by Mark, Malte, Rebecca et al. FCAL is 'green' for the run. A test stand has been set up with 60 bases in EEL118, that will undergo significant testing protocols, including measuring voltages (board probe tests). Three strands of consecutive boards have had new firmware uploaded, which it is hoped will be less susceptible to locking when the beam is on; these are easy to spot on the occupancy plots and a note will be made for shift takers to keep an eye on them. Matt reported that work has been done on the remote programming of the bases and understanding failure modes and Paul Smith contributed to this effort.
    2. BCAL: BCAL is in the 'green'. There is an issue on missing tracks in the detector: possible suspects are the Start Counter and maybe the BCAL. There was an issue with BCAL LV modules being off, but this was fixed a week ago; the tracking issue persists in runs since then. BCAL timing was discussed in connection to cosmics runs. Extensive cosmics for tracking are not taken frequently, so perhaps these issues need more study.
    3. CCAL: no report.
  3. Calibration: Sean proposed that we start the run with the same calibrations from last time and then adjust as needed.
  4. Monitoring
    1. BCAL LED: Varun has taken 5-deg-chiller runs for the SiPM damage study, Mark and Nick will change the chiller temps fo 10 deg and 18 deg, one day each, and Varun will take that data. The chillers will return to 5 deg for the run.
    2. BCAL (offline): Zisis will get going on this soon and work with Tolga.
    3. FCAL (offline): Susan reported that when she tried to correct data fluctuations by dividing with the LED efficiency this resulted in an over correction. One of the issues is that some sub-runs have only data and no LED events. Mark mentioned that he will continue with the nominal 6 min pulsing cycle (LED colours, LV/HV and off). Susan will look at the automatically skimmed LED events as soon we start running. She will also look at past runs where the LED pulsing frequency was larger; some runs had 1kHz and PMT/base drooping was seen (see log book entry 3595928).
  5. Reconstruction: no report.
  6. Simulations: no report.
  7. Action Items
    1. Short Term: Item #1 was discussed: Sean proposed that the 2020 constants be put into the databases and their verification can be a longer term item. Susan volunteered to help Mark with consistency checks.
    2. Long Term: Tegan started on Item #2.
    3. Work Packages: no updates.
  8. Any other business: none.